OpenStudy (anonymous):

So, our professor just returned us our test, and I got all questions correct except for one: "Solve for \(x\):\[24=16x-40"\]I don't get it: I did everything right! This is my work, step by step:\[24=16x-40,\]\[24+40=16x,\]\[64=16x,\]\[x=\frac{64}{16},\]\[x=\frac{\not{6}4}{1\not{6}}=\frac{4}{1}=4.\]What do you guys think!?

OpenStudy (4n1m0s1ty):

its right but for the wrong reason

OpenStudy (ash2326):

It's correct. Wonder why did he mark it incorrect

OpenStudy (4n1m0s1ty):

last line should be 64/16

OpenStudy (4n1m0s1ty):

you did 6*4/1*6 = 4

OpenStudy (4n1m0s1ty):

Minor mistake though to me. I wouldn't have counted off the entire problem

OpenStudy (anonymous):

the answer is right! you should tell him and get the grade you deserve. and I was wondering if that was possible. the 64/16 and reducing the 6's I was amazed by the technique at first. @4n1m0s1ty that that he did is incorrect? as in not possible?

OpenStudy (anonymous):

LOl. I bet he must have marked the problem wrong seeing the last step

OpenStudy (anonymous):

He also said that \(1=\sqrt{1}=\sqrt{(-1)(-1)}=\sqrt{-1}\sqrt{-1}=i\cdot i=-1\).

OpenStudy (anonymous):

The above statement is wrong

OpenStudy (anonymous):

You can never take product of square root of two negative numbers

OpenStudy (4n1m0s1ty):

yeah thats not a legitimate operation if i did 15/5 --> 1 its not the same

OpenStudy (anonymous):

Maybe getting back at him with a mathematical fallacy was not a good idea. :P

OpenStudy (anonymous):

true @4n1m0s1ty

OpenStudy (anonymous):

LOL. Make sure you never cancel the way you did in the last step. You can cancel in the following way